OpenStudy (anonymous):

Let y = y1(x) be a solution of (H): y'' + p(x)y' + q(x)y = 0 where p and q are continuous function on an interval I. Let a be in I and assume that y1(x) does not =0 on I. Set y2(x) = y1(x) integral [(e^-integral p(u)du)/(y1^2(t))] dt. Show that y2 is a solution of (H) and that y1 and y2 are linearly independent. Show that

OpenStudy (anonymous):

Would you by chance be able to help with this?? @zepdrix

zepdrix (zepdrix):

Hmm I vaguely remember that formula. Reduction of Order I believe it's called. Used for finding a second solution when you already have a first. Hmmmmm...

OpenStudy (anonymous):

I have not heard of that... I will google, but I am at a total lose on what to do

OpenStudy (sirm3d):

suppose there exists some function \(v(t)\) such that \(y(t)=v(t) y_1(t)\) is a solution to the DE. This is the reduction of order method.

OpenStudy (sirm3d):

put \(y=vy_1\) \[\large{y'=v'y_1+vy_1'\\y''=v''y_1+2v'y_1'+vy_1''\\0=y''+py'+qy=(v''+2v'y_1'+vy_1'')+p(v'y_1+vy_1')+q(vy_1)\\=v''+2v'y_1'+pv'y_1+v(y_1''+py_1'+qy_1)\\=v''+(2y_1'+py_1)v'+0}\]

OpenStudy (anonymous):

so where is y2? or are you saying that y is y2?